It took a boat 2 hours to rich town A going upstream. The way back was one hour and 20 minutes. What is the speed of the boat in still water if the speed of the stream is 3mph?
Answer by Alan3354(69443)

Same problem, different numbers.
=========================================
It took a plane 2 hours to make a trip with the head wind and 1.5 hours to get back with the tail wind. If the speed of the wind was 30mph, how far apart are the two cities?
-----------
a = plane's airspeed
---
d = r*t = (a-30)*2 = (a+30)*1.5
2a - 60 = 1.5a + 45
a = 210 mi/hr
d = (210-30)*2
d = 360 miles